OpenStudy (hihi):

The formula for the circumference C of a circle with radius r is C = 2πr. Which of the following is a formula for π (pi) of the circle in terms of the circumference and radius? A. PI=C+2r B: C-2r C: PI=2R/C D: PI/C/2R

OpenStudy (anonymous):

\[C=2\pi*r\] Solve for \[\pi\] \[\pi=C/(2r)\] Choice D

jhonyy9 (jhonyy9):

D. is right